### v1

@@ 5.3 Visual diagnostics
We now present three diagnostic tools that help clarify the spatial confounding adjustments made by spatial regression models, as well as the composition of the causal contrast they build across the spatial dataset. Our proposed diagnostics build on the weighting diagnostics introduced by Chattopadhyay and Zubizarreta (2023). We apply these diagnostics to the Superfund sites and birth outcomes dataset, examining each of the three spatial regression models described in Section 3.
The first two diagnostics visually characterize the class of unmeasured confounders that spatial regression adjusts for. Figure 2 illustrates the first diagnostic tool, which plots the maximum absolute conditional bias arising from $\bm{U}$ against its Moran’s I statistic and its coefficient in the outcome model. For this diagnostic, we employ nonlinear programming because the analytical bound in Proposition 5.1 can sometimes be quite conservative. For all three spatial regression models, this diagnostic reveals that the maximum absolute conditional bias decreases as the unmeasured confounder’s Moran’s I increases, and increases with the magnitude of $\gamma$, the coefficient of $\bm{U}$ in the outcome model. This pattern empirically corroborates Proposition 5.1.
Figure 2: Maximum absolute conditional bias as a function of Moran’s I and absolute regression coefficient $|\gamma|$.
@@
The second diagnostic, illustrated in Figure 7 of the Supplementary Material, displays three instances of $\bm{U}$ together with their respective Moran’s I statistics and mean imbalances. As expected, confounders characterized by higher Moran’s I tend to exhibit reduced mean imbalance. In particular, the three spatial regression models target unmeasured confounders with different forms of spatial autocorrelation. For the random effects model, spatially autocorrelated unmeasured confounders are nearly constant within states. In contrast, the conditional autoregressive model yields spatially autocorrelated unmeasured confounders that are approximately constant over adjacencies; here, sites $i$ and $j$ are defined as adjacent if site $j$ is one of site $i$’s five nearest neighbors or vice versa. For the Gaussian process model, spatially autocorrelated unmeasured confounders appear to be smooth, continuous functions of spatial coordinates.
Figure 3: Implied weights $w_{1},\ldots,w_{n}$ in three common spatial regression models.
@@
The third diagnostic, shown in Figure 3, displays the implied weights $w_{1},\ldots,w_{n}$ from the three spatial regression models. For the random effects model, the weights are largest in magnitude for Superfund sites located in states with a high proportion of control sites, or for control sites in states with a high proportion of treated sites. For instance, sites in Idaho and Montana, which contain only controls, are assigned weights that are nearly zero. For the conditional autoregressive model, the largest weights correspond to treated sites bordering many control sites, or vice versa. For the Gaussian process model, the largest weights are observed for Superfund sites that are geographically proximate to sites of the opposite treatment group.
@@
In sum, Figure 3 shows that spatial regression produces a weighting estimator in which the treated and control components are weighted to be spatially proximate, according to a definition of proximity that is model-specific. Further discussion of this spatially localized weighting is presented in Supplementary Section 11.5. This phenomenon is akin to distance-adjusted propensity score matching, which pairs treated and control units with similar propensity scores only if they are geographically close (Papadogeorgou et al. 2019) and geographic regression discontinuity designs, which construct treated-control contrasts using observations situated on either side of a geographic boundary separating treatment groups (Keele et al. 2015). However, unlike these methods, spatial regression imposes a very strict requirement of linearity, and if there are any deviations from a simple linear outcome model, its estimates can be subject to large biases. See Supplementary Section 11.6 for examples. In the next section, we present an extension of spatial regression that is capable of addressing nonlinearities and treatment effect heterogeneity in the outcome model.
6 Extending spatial regression estimators for interpretability and robustness
@@
In this section, we extend spatial regression to estimate the average treatment effect on the treated, introducing a weighting estimator that adjusts for multiple forms of unmeasured spatial confounding, relaxes the assumptions of linearity and effect homogeneity, and is constructed to be sample bounded.
@@
Let $X^{\text{aug}}=(\bm{X},\bm{V})\in\mathbb{R}^{n\times(12+J)}$ be the set of measured covariates, augmented with $J$ pre-selected eigenvectors $\bm{V}=(\bm{v}_{1},\ldots,\bm{v}_{J})\in\mathbb{R}^{n\times J}$ of any positive semidefinite matrix $\bm{S}$. Building on Wang and Zubizarreta (2020) and Zubizarreta (2015), we solve the following quadratic programming problem in the control group:
@@
The solution to this optimization problem yields a set of normalized, minimum-variance control weights that approximately balance the basis functions $B(\cdot)=\{B_{1}(\cdot),\ldots,B_{K}(\cdot)\}^{T}$ relative to the treated group. To ensure that the resulting estimator is sample bounded (Robins et al. 2007), we also impose non-negativity constraints on the weights. The spatial weighting estimator is then defined as $\widehat{\tau}_{SW}=\frac{1}{n_{t}}\sum_{i:Z_{i}=1}Y_{i}-\sum_{i:Z_{i}=0}w_{i}Y_{i},$ where $\{w_{i}\}_{Z_{i}=0}$ are the optimized weights obtained from the mathematical program described above.
@@
Our weighting approach requires selecting $J$ augmenting eigenvectors $\bm{v}_{1},\ldots,\bm{v}_{J}$. To reduce bias from the main classes of unmeasured spatial confounding implicitly targeted by the random effects, conditional autoregressive, and Gaussian process models, we include a large number of eigenvectors corresponding to the highest eigenvalues from each model. Unlike methods that adjust for only a single confounder class, $\widehat{\tau}_{SW}$ enables simultaneous adjustment for multiple types of unmeasured confounders, including those constant at the state level, those exhibiting smooth spatial variation, and those with minimal variation across adjacencies. Proposition 11.3 further characterizes the finite-sample bias and variance of the spatial weighting estimator as a function of the Moran’s I statistic of the unmeasured spatial confounder. In the next section, we assess the performance of the spatial weighting estimator in a simulation study.
7 Simulation-based analysis of forms of adjustment
For $i=1,\ldots,n=1583,$ we generate unmeasured confounders $\bm{U}$ using three different mechanisms, each inducing a different spatial autocorrelation structure. For all three mechanisms, we generate $\bm{U}$ with a two-stage procedure: 1) we simulate $U_{i}\sim\mathcal{N}(Z_{i},0.1^{2})$ and 2) we set $U_{i}=\sum_{j=1}^{n}W_{ij}U_{j}/\sum_{j=1}^{n}W_{ij}$. The first class of unmeasured confounders (cluster-based) is designed to be constant within states, hence we define $W_{ij}=I(C_{i}=C_{j})$. The second class of unmeasured confounders (adjacency-based) is designed to be smooth across adjacencies, hence we define $W_{ij}=(A^{100})_{ij}$. The third class of unmeasured confounders (distance-based) is designed to vary smoothly with geographical distance, hence we define $W_{ij}=e^{-d_{ij}/50\text{km}}$. Supplementary Section 11.8 provides a plot of one observation of $\bm{U}$ from each class.
We further consider three possible outcome models. The linear outcome model generates outcome as $Y_{i}=\bm{\beta}^{T}\bm{X}_{i}+Z_{i}-0.2U_{i}+\mathcal{N}(0,0.1^{2})$. The linear-interaction outcome model generates outcome as $Y_{i}=\bm{\beta}^{T}\bm{X}_{i}+Z_{i}-0.5U_{i}+U_{i}Z_{i}+\mathcal{N}(0,0.1^{2})$. The nonlinear-interaction outcome model generates outcome as $Y_{i}=\bm{\beta}^{T}\bm{X}_{i}+\tau Z_{i}+Z_{i}\sin(U_{i})-U_{i}^{2}+U_{i}Z_{i}(X_{2i}+1)+0.3Z_{i}X_{1i}^{2}+\mathcal{N}(0,0.1^{2})$. For each of the $9$ data-generating scenarios produced by the three classes of unmeasured confounders and three outcome models, we create $M=500$ datasets of size $n=1583$, with $\bm{X}$ and $\bm{Z}$ drawn directly from our Superfund dataset.
Under each data-generating scenario, we estimate the ATT using six different approaches. The first approach (OLS) fits an ordinary least squares regression of $Y$ on $X$ and $Z$ and extracts the coefficient estimate of $Z$. Approaches two through four employ generalized least squares to fit the random effects (RE), conditional autoregressive (CAR), and Gaussian process (GP) models described in Section 3, likewise extracting the coefficient estimate of $Z$. The fifth approach (spatial coordinates) incorporates the latitude and longitude of each Superfund site as two additional covariates in an augmented inverse propensity score estimator, following the framework proposed by Gilbert et al. (2021). The sixth approach (spatial weighting) implements our proposed methodology from Section 6. In order to adjust for a heterogeneous class of unmeasured confounders with spatial autocorrelation, we augment $\bm{X}$ with $50$ eigenvectors corresponding to the largest eigenvalues from each spatial regression model above, resulting in $J=150$. We use the package sbw for implementation (Zubizarreta 2015).
We evaluate the performance of the six approaches in estimating the ATT through bias and root mean squared error. Figure 4 presents boxplots of the ATT estimates and Supplementary Section 11.8 presents numerical results.
We briefly highlight two key observations. First, the results under the linear outcome model, in the first row of the boxplot grid, provide empirical support for our findings in Section 5. Specifically, each spatial regression model yields unbiased estimates for the class of unmeasured confounders it targets: the random effects model under cluster-based confounding; the conditional autoregressive model under adjacency-based confounding; and both the conditional autoregressive and Gaussian process models under distance-based confounding. However, when nonlinearity or treatment effect heterogeneity by $\bm{U}$ is present in the outcome model, all three estimators suffer from bias.
Adjustment for spatial coordinates can reduce bias from unmeasured spatial confounding when the confounder is a measurable, or nearly continuous, function of spatial coordinates and the exposure exhibits non-spatial variation (Gilbert et al. 2021). These assumptions are not explicitly satisfied by any of the confounding mechanisms considered, and the spatial coordinates estimator exhibits substantial bias across most data-generating scenarios.
@@ 7 Simulation-based analysis of forms of adjustment
In contrast, our proposed spatial weighting estimator remains approximately unbiased across all nine data-generating scenarios and — in terms of bias — outperforms each competing model in seven of the nine scenarios, although sometimes exhibits larger uncertainty (see Supplementary Section 11.8). These simulation results demonstrate that the spatial weighting estimator can flexibly accommodate complex nonlinearities and treatment effect heterogeneity, while simultaneously adjusting for unmeasured confounders that are constant within states, smooth across adjacencies, or vary continuously over space.
@@ 8 Effects of Superfund remediation on birth outcomes
We now apply our proposed spatial weighting approach to estimate the effect of Superfund site remediation during 2001–2015 on five birth outcomes measured during 2016–2020 in New Jersey $(n=139)$. All five birth outcome variables were obtained at the municipal-level from the New Jersey State Health Assessment and subsequently aggregated to the 2 km buffer level around each Superfund site.
The first outcome is the percent of low birth weight (birth weight $\leq 2500$ grams) births; the second outcome is average birth weight in pounds; the third outcome is infant mortality (death $\leq 1$ year of age) rate; the fourth outcome is preterm birth rate (gestational age <37<37 weeks); and the fifth outcome is average gestational age. These publicly available, population-wide data cover all births, regardless of insurance status or hospital affiliation, at a fine spatial resolution (across $564$ municipalities). New Jersey contains the highest number of Superfund sites out of all US states. Our target of estimation is the causal risk ratio, defined as $\mathbb{E}(Y(1)|Z=1)/\mathbb{E}(Y(0)|Z=1)=\mathbb{E}(Y|Z=1)/(\mathbb{E}(Y|Z=1)-\tau_{ATT})$. Standard errors are derived from bootstrapping with $500$ replications. Figure 10 plots the estimated causal risk ratio as the number of hidden covariates ($2J$) incorporated into the spatial weighting method increases. For $J\in\{1,\ldots,20\}$, we append a total of $J$ highest-eigenvalue eigenvectors of $\bm{S}^{\text{CAR}}$ and $J$ highest-eigenvalue eigenvectors of $\bm{S}^{\text{GP}}$ to the design matrix. The point estimates, although accompanied with large uncertainty, tend to diverge from the null until $J=9$ and then plateau. Since the stabilization occurs at $9$, we adopt $J=9$ for a total of $18$ hidden covariates in the final specification of the spatial weighting estimator.
Table 1 reports estimated causal risk ratios for each of the five birth outcomes under each of the six modeling approaches. Point estimates are broadly similar across the six models, suggesting that unmeasured spatial confounding bias may be modest. The spatial weighting approach yields the lowest estimates for low birth weight and infant mortality — 0.93 (95 % CI 0.86, 1.02) and 0.67 (95 % CI 0.50, 1.01), respectively — with smaller uncertainty than the other four spatial models.
Superfund site remediation may improve birth outcomes by reducing maternal exposure to hazardous substances such as heavy metals, solvents, and air- or waterborne contaminants — such as arsenic, lead, and PFAS — known to affect fetal development (Henn et al. 2016; Keeler et al. 2023). Cleanup efforts may also lead to secondary improvements in neighborhood conditions (e.g., reduced stress, improved housing or infrastructure, neighborhood investment, increased housing prices) that may indirectly improve birth outcome measures (Currie 2011; Kiel and Zabel 2001; Noonan et al. 2007).
Our findings are somewhat consistent with prior research. Currie et al. (2011) report that remediation may reduce the incidence of congenital anomalies by 20–25%, and also find a modest reduction in preterm birth, although no significant effects on low birth weight or infant mortality. In related work, Currie (2011) documents a demographic shift in the population of mothers residing near cleaned sites, with increases in the proportion of White and college-educated mothers, perhaps due to housing price appreciation. Such demographic shifts may contribute to the estimated effects observed in our study. However, the existing literature is limited; most prior research has focused on associations between residential proximity to Superfund sites and birth outcomes, rather than on estimating the causal impact of site remediation (Brender et al. 2011; Henn et al. 2016; Kihal-Talantikite et al. 2017; Langlois et al. 2009).
Rate of Low Birth Weight Average Birth Weight Infant Mortality Rate Preterm Birth Rate Average Gestational Age
OLS 0.96 (0.90, 1.03) 1.00 (1.00, 1.01) 0.83 (0.63, 1.21) 0.96 (0.90, 1.03) 1.00 (1.00, 1.00)
RE 0.96 (0.87, 1.07) 1.00 (1.00, 1.01) 0.80 (0.56, 1.42) 0.96 (0.87, 1.06) 1.00 (1.00, 1.00)
CAR 0.94 (0.85, 1.05) 1.00 (0.99, 1.01) 0.78 (0.54, 1.36) 0.96 (0.88, 1.06) 1.00 (1.00, 1.00)
GP 0.96 (0.87, 1.07) 1.00 (0.99, 1.01) 0.81 (0.57, 1.37) 0.96 (0.88, 1.07) 1.00 (1.00, 1.00)
SC 0.96 (0.87, 1.06) 1.00 (0.99, 1.01) 0.81 (0.59, 1.28) 0.96 (0.88, 1.05) 1.00 (1.00, 1.00)
SW 0.93 (0.86, 1.02) 1.00 (0.99, 1.00) 0.67 (0.50, 1.01) 0.96 (0.88, 1.07) 1.00 (1.00, 1.00)
Table 1: Estimates of the causal risk ratio and 95% confidence intervals from six methods, assessing the impact of Superfund site remediation (2001–2015) on five birth outcomes (2016–2020) within 2 kilometers of Superfund sites in New Jersey. We define the causal risk ratio as $\mathbb{E}(Y(1)|Z=1)/\mathbb{E}(Y(0)|Z=1)$. The five birth outcomes are: 1) average birth weight (pounds), 2) gestational age (weeks), 3) infant mortality rate (%), 4) rate of low birth weight births (%), and 5) preterm birth rate (%). OLS, ordinary least squares; RE, random effects; CAR, conditional autoregressive; GP, Gaussian process; SC, spatial coordinates; SW, spatial weighting.
Figure 5 presents diagnostic plots summarizing the spatial weighting procedure: panel (a) shows covariate balance, including the 18 hidden spatial covariates; panel (b) displays a histogram of the weights; panel (c) maps the weights in geographic space; and panel (d) plots each of the hidden spatial covariates. Approximate balance on the 18 hidden spatial covariates yields a weighted contrast between treated and control units that are geographically close. Although the effective sample size is only $(\sum_{i=1}^{n}w_{i}^{2})^{-1}=45.9$, limiting statistical power, our analysis demonstrates the utility of spatial weighting in a challenging setting. Specifically, it highlights the method’s interpretability, its ability to incorporate spatial information via hidden spatial covariates, and its potential for transparent diagnostic evaluation.
Figure 5:
@@
population density 2504.165 3697.645 3697.645
proportion Hispanic 0.097 0.075 0.075
proportion Black 0.110 0.115 0.115
proportion Indigenous 0.006 0.006 0.006
proportion Asian 0.052 0.046 0.046
median household income 59855.855 62735.665 62735.665
median house value 164254.979 183720.081 183720.081
proportion in poverty 0.073 0.056 0.056
proportion graduated HS 0.817 0.852 0.852
median year built 1966.374 1962.972 1962.972
Site Score 45.654 42.085 42.085
$\bm{v}_{1}^{\text{CAR}}$ 0.003 -0.019 -0.017
$\bm{v}_{2}^{\text{CAR}}$ -0.003 0.018 0.021
$\bm{v}_{3}^{\text{CAR}}$ 0.002 -0.009 -0.012
$\bm{v}_{4}^{\text{CAR}}$ 0.003 -0.015 -0.018
$\bm{v}_{5}^{\text{CAR}}$ -0.003 0.015 0.018
$\bm{v}_{6}^{\text{CAR}}$ 0.003 -0.015 -0.020
$\bm{v}_{7}^{\text{CAR}}$ -0.004 0.023 0.024
$\bm{v}_{8}^{\text{CAR}}$ 0.001 -0.004 -0.010
$\bm{v}_{9}^{\text{CAR}}$ 0.004 -0.016 -0.024
$\bm{v}_{1}^{\text{GP}}$ -0.085 -0.085 -0.085
$\bm{v}_{2}^{\text{GP}}$ 0.003 -0.022 -0.021
$\bm{v}_{3}^{\text{GP}}$ 0.000 -0.002 0.002
$\bm{v}_{4}^{\text{GP}}$ 0.001 -0.004 -0.004
$\bm{v}_{5}^{\text{GP}}$ 0.001 -0.011 -0.007
$\bm{v}_{6}^{\text{GP}}$ -0.005 0.030 0.037
$\bm{v}_{7}^{\text{GP}}$ 0.006 -0.030 -0.041
$\bm{v}_{8}^{\text{GP}}$ 0.001 0.003 -0.005
$\bm{v}_{9}^{\text{GP}}$ -0.002 0.015 0.013
(a) Balance and representativeness.

### v2

@@ 5.3 Visual diagnostics
We now present three visual diagnostics that help clarify the spatial confounding adjustments made by spatial regression models, as well as the composition of the causal contrast they build across the spatial dataset. Our visual diagnostics build on the weighting diagnostics introduced by Chattopadhyay and Zubizarreta (2023). We apply them to the Superfund sites and birth outcomes dataset, examining each of the three spatial regression models described in Section 3.
@@
The first two diagnostics depict the class of unmeasured confounders that spatial regression adjusts for. Figure D in the Supplementary Material illustrates the first visual diagnostic, which plots the maximum absolute conditional bias arising from $\bm{U}$ against its Moran’s I statistic and its coefficient in the outcome model. For this diagnostic, we employ nonlinear programming because the analytical bound in Proposition 5.1 can sometimes be quite conservative. For all three spatial regression models, this diagnostic reveals that the maximum absolute conditional bias decreases as the unmeasured confounder’s Moran’s I increases, and increases with the magnitude of $\gamma$, the coefficient of $\bm{U}$ in the outcome model. This pattern empirically corroborates Proposition 5.1. The second visual diagnostic, illustrated in Figure 7 of the Supplementary Material, displays three instances of $\bm{U}$ together with their respective Moran’s I statistics and mean imbalances. As expected, confounders characterized by higher Moran’s I exhibit reduced mean imbalance. However, the three spatial regression models target unmeasured confounders with different forms of spatial autocorrelation. For the random effects model, spatially autocorrelated unmeasured confounders are nearly constant within states. The conditional autoregressive model targets spatially autocorrelated unmeasured confounders that are approximately constant over adjacencies; here, sites $i$ and $j$ are defined as adjacent if site $j$ is one of site $i$’s five nearest neighbors or vice versa. For the Gaussian process model, spatially autocorrelated unmeasured confounders appear to be smooth functions of spatial coordinates.
Figure 2: Implied weights $w_{1},\ldots,w_{n}$ in three common spatial regression models. Our analysis enables a direct comparison of the spatial adjustments each model applies. In Florida, the random effects model produces weights that are nearly uniform across the control sites in the state. The Gaussian process model assigns higher weights to control sites that are geographically closer to treated sites, leading to greater variation in control weights. The conditional autoregressive model places the largest weights on control sites adjacent to treated sites, and some control sites receive negative weights.
@@
The third diagnostic, shown in Figure 2, displays the implied weights $w_{1},\ldots,w_{n}$ from the three spatial regression models. For the random effects model, the weights are largest in magnitude for Superfund sites located in states with a high proportion of control sites, or for control sites in states with a high proportion of treated sites. For instance, sites in Montana and Wyoming, which contain only control sites, are assigned weights that are nearly zero. For the conditional autoregressive model, the largest weights correspond to treated sites bordering many control sites, or vice versa. For the Gaussian process model, the largest weights are observed for Superfund sites that are geographically proximate to sites of the opposite treatment status. Across all three models, some control sites receive negative weights, producing an estimator that is not sample bounded (Robins et al. 2007).
@@
In summary, Figure 2 shows that spatial regression produces a weighting estimator in which the treated and control components are weighted to be spatially proximate, according to a model-specific definition of proximity. Further discussion of this spatially localized weighting is presented in Supplementary Section E. This approach is akin to distance-adjusted propensity score matching, which pairs treated and control units with similar propensity scores only if they are geographically close (Papadogeorgou et al. 2019) and geographic regression discontinuity designs, which construct treated-control contrasts using observations situated on either side of a geographic boundary separating treatment groups (Keele et al. 2015). However, unlike these methods, spatial regression imposes a very strict requirement of linearity, and if there are deviations from this specification, its estimates can be subject to biases. See Supplementary Section F for examples. In the next section, we present an extension of spatial regression that is capable of addressing nonlinearities and heterogeneity of treatment effects in the outcome model.
6 Extending spatial regression for robustness
@@
In this section, we extend spatial regression to estimate the average treatment effect on the treated, introducing a weighting estimator that adjusts for multiple forms of unmeasured spatial confounding, relaxes the assumptions of linearity and effect homogeneity, and can be constructed to be sample bounded.
@@
Let $X^{\text{aug}}=(\bm{X},\bm{V})\in\mathbb{R}^{n\times(18+J)}$ be the set of measured covariates, augmented with $J$ pre-selected eigenvectors $\bm{V}=(\bm{v}_{1},\ldots,\bm{v}_{J})\in\mathbb{R}^{n\times J}$ of any positive semidefinite matrix $\bm{S}$. Building on Wang and Zubizarreta (2020) and Zubizarreta (2015), we solve the following quadratic programming problem in the control group:
@@
The solution to this optimization problem yields a set of normalized, minimum-variance control weights that approximately balance the basis functions $B(\cdot)=\{B_{1}(\cdot),\ldots,B_{K}(\cdot)\}^{T}$ relative to the treated group. To ensure that the resulting estimator is sample bounded (Robins et al. 2007), we also impose non-negativity constraints on the weights. The spatial weighting estimator is then defined as $\widehat{\tau}_{SW}=(n_{t})^{-1}\sum_{i:Z_{i}=1}Y_{i}-\sum_{i:Z_{i}=0}w_{i}Y_{i},$ where $\{w_{i}\}_{Z_{i}=0}$ are the optimized weights obtained from the mathematical program described above.
@@
Our weighting approach requires selecting $J$ augmenting eigenvectors $\bm{v}_{1},\ldots,\bm{v}_{J}$. To reduce bias from the main classes of unmeasured spatial confounding implicitly targeted by the random effects, conditional autoregressive, and Gaussian process models, we include a large number of eigenvectors corresponding to the highest eigenvalues from each model. Unlike methods that adjust for only a single confounder class, $\widehat{\tau}_{SW}$ enables simultaneous adjustment for multiple types of unmeasured confounders, including those constant at the state level, those exhibiting smooth spatial variation, and those with minimal variation across adjacencies. Proposition G.1 characterizes the finite-sample bias and variance of the spatial weighting estimator as a function of the Moran’s I statistic of the unmeasured spatial confounder.
@@
In Supplementary Section H, we assess the performance of the spatial weighting estimator in simulation. First, we investigate whether it can mitigate confounding bias from three distinct classes of unmeasured confounders with different patterns of spatial autocorrelation. Second, we evaluate its performance while varying the complexity of the outcome model. The simulation results demonstrate that the spatial weighting estimator can flexibly accommodate complex nonlinearities and heterogeneity of treatment effects, while simultaneously adjusting for unmeasured confounders that are constant within states, smooth across adjacencies, or vary continuously over space.
@@ 7 Effects of Superfund remediation on birth outcomes
We now apply our spatial weighting approach to estimate the effect of Superfund site remediation during 1991–2015 on the rates of congenital anomalies and small vulnerable newborns during 2016–2018 among Medicaid-covered births in the United States. During this period, there were approximately $11.6$ million total U.S. births, of which 4,361,860 (37%) were Medicaid-covered births identified in the inpatient claims data available to us. Of these, 266,604 occurred within Superfund site areas or their 2-kilometer buffers. For stable estimation, we excluded sites with fewer than $10$ Medicaid-covered live births within the site area and its 2-kilometer buffer, resulting in a final dataset of $n=1079$ Superfund sites (see Supplementary Section I for additional details). To our knowledge, this study represents the first national-scale linkage of Medicaid claims with EPA remediation data to evaluate health impacts of Superfund cleanup.
We report results for five methods: OLS, three spatial regression models (RE, CAR, GP), and the proposed spatial weighting (SW) approach, each with confidence intervals obtained via bootstrap with 500 replications. In the final specification of the spatial weighting estimator, we add the $10$ highest-eigenvalue eigenvectors from each of $\bm{S}^{\text{RE}}$, $\bm{S}^{\text{CAR}}$ and $\bm{S}^{\text{GP}}$ to the design matrix. We apply balancing thresholds $\delta_{k}$ of $0.001$ to the $17$ measured covariates and $0.01$ to the $30$ latent covariates after standardization.
Remediation appears to have a modest beneficial effect on the percentage of small vulnerable newborns, with ATT estimates ranging from 0.3-0.3% to 0.8-0.8% (relative to an overall rate of 14.2% among Medicaid-covered births), but no discernible effect on the percentage of congenital anomalies. The spatial weighting approach yields ATT estimates of 0.60-0.60% (95% CI: 1.15-1.15%, 0.10-0.10%) for the rate of small vulnerable newborns and 0.070.07% (95% CI: 0.72-0.72%, 0.910.91%) for the rate of congenital anomalies, both with smaller confidence intervals than the other three spatial models. Estimates for all methods are reported in Table 1.
Figure 3: Estimate of the ATT $\mathbb{E}(Y(1)-Y(0)\mid Z=1)$ and effective sample size (ESS) $(\sum_{i=1}^{n}|w_{i}|)^{2}/\sum_{i=1}^{n}w_{i}^{2}$ obtained from the spatial weighting approach as a function of (a) the number of highest-eigenvalue eigenvectors from $\bm{S}^{\text{RE}}$, $\bm{S}^{\text{CAR}}$, and $\bm{S}^{\text{GP}}$ included as latent covariates, fixing the balancing threshold to 0.01, and (b) the balancing threshold applied to the latent covariates after standardization, fixing the number of latent covariates included from each model to $10$ ($30$ in total).
Adjusting for unmeasured spatial confounding, whether through spatial regression adjustment or our proposed weighting approach, attenuates the estimated effects of remediation on the percentage of small vulnerable newborns. This attenuation can be observed in (i) Table 1, where spatial weighting and spatial regression yield estimates closer to the null than OLS; (ii) Figure 3(a), where including a larger number of latent covariates in the spatial weighting approach attenuates the ATT (blue) while reducing the effective sample size (green); and (iii) Figure 3(b), where enforcing tighter balance on the 30 latent covariates similarly attenuates the ATT (blue) while reducing the effective sample size (green).
% Small Vulnerable % with Congenital Anomalies
Ordinary least squares -0.74 (-1.24, -0.25) 0.00 (-0.73, 0.80)
Random effects -0.52 (-1.37, 0.04) -0.30 (-1.35, 1.09)
Conditional autoregressive -0.42 (-1.34, 0.12) 0.29 (-1.01, 1.50)
Gaussian process -0.36 (-1.53, 0.23) -0.058 (-1.31, 1.12)
Spatial weighting -0.60 (-1.15, -0.10) 0.07 (-0.72, 0.91)
Table 1: Estimates of the ATT and 95% confidence intervals from five methods, assessing the impact of Superfund site remediation (1991–2015) on two birth outcomes (2016–2018) among Medicaid-covered births within 2 kilometers of Superfund site areas. The two birth outcomes are: i) the percentage of newborns that are “small vulnerable”, meaning preterm, low birth weight, or small for gestational age, and ii) percentage of newborns with congenital anomalies.
Figure 4 demonstrates how spatial weighting approximates a hypothetical experiment across space, providing visual diagnostics that link geography, confounding adjustment, and effect estimation. Panels (a) and (b) present covariate means within each treatment group before and after spatial weighting for the 17 measured and 30 latent covariates, respectively. The control group is reweighted so that its covariate means closely align with those of the treated group. The balance tables illustrate how the spatial weighting procedure achieves the specified balance constraints, giving the investigator explicit control over which covariates are balanced and to what degree.
Panel (c) displays a histogram of the weights. The effective sample size for the spatial weighting approach is $539.9$, compared to $597.3$, $577.6$, $563.8$, and $569.4$ for OLS, RE, CAR, and GP, respectively; the nominal sample size is $n=1079$. Although the effective sample size for the spatial weighting approach is lower than for the three spatial regression models, the confidence interval is in fact narrower, suggesting that approximately balancing the observed covariates and simultaneously adjusting for three types of latent covariates may reduce variance by explaining residual spatial variation in outcome that is independent of observed covariates and treatment (see Supplementary Section G and Figure 12).
Panel (d) plots the 30 latent covariates. This diagnostic reveals the latent covariates that spatial regression models balance implicitly and that the spatial weighting approach balances explicitly. By approximately balancing this set of latent covariates, the spatial weighting approach simultaneously adjusts for three classes of unmeasured spatial confounders: those that are constant within states, those that exhibit neighborhood-level similarity, and those that vary smoothly over space.
Panel (e) maps the weights in geographic space. By balancing on the 30 latent covariates, the spatial weighting approach constructs a geographically local contrast between treated and control sites. A substantial share of this contrast (17.7% of the total weight) arises from the New York–New Jersey region, reflecting the high concentration of both treated and control Superfund sites there. There are 114 control sites that receive zero weight, with many located far from any treated site. Although spatial weighting reduces the effective sample size, it ensures that inference is restricted to treated and control sites that are comparable: control sites far from treated sites may differ with respect to unmeasured spatial confounders and should therefore be excluded from the comparison.
Figure 4: Diagnostic plots summarizing the spatial weighting procedure.
@@
Population density (ppl/mi2) 1730 1610 1610
Proportion of Hispanic residents 0.0681 0.0548 0.0549
Proportion of Black residents 0.104 0.118 0.118
Proportion of Asian residents 0.0238 0.0189 0.0189
Proportion of Indigenous residents 0.00814 0.00821 0.00818
Proportion of renter occupied housing 0.325 0.296 0.295
Median household income (\$) 32,500 31,700 31,700
Median house value (\$) 94,400 87,100 87,000
Proportion of residents in poverty 0.115 0.124 0.124
Proportion of residents that graduated high school 0.742 0.737 0.737
Median year of housing construction 1960 1960 1960
Site score 43.7 38.4 38.4
Metal present 0.744 0.753 0.754
VOC present 0.751 0.534 0.533
Contamination through soil 0.845 0.816 0.815
Contamination through liquid medium 0.896 0.800 0.800
Contamination through gas 0.145 0.0670 0.0667
(a) Balance and representativeness: measured covariates.
Control Treated
Covariate Before weighting After weighting
$\bm{v}_{1}^{\text{RE}}$ -0.00887 -0.00786 -0.00815
$\bm{v}_{2}^{\text{RE}}$ -0.00775 -0.01100 -0.01130
$\bm{v}_{3}^{\text{RE}}$ -0.00858 -0.00609 -0.00639
$\bm{v}_{4}^{\text{RE}}$ -0.00909 -0.00427 -0.00406
$\bm{v}_{5}^{\text{RE}}$ -0.00745 -0.00625 -0.00596
$\bm{v}_{6}^{\text{RE}}$ -0.00648 -0.00972 -0.01000
$\bm{v}_{7}^{\text{RE}}$ -0.00596 -0.01010 -0.00996
$\bm{v}_{8}^{\text{RE}}$ -0.00673 -0.00343 -0.00313
$\bm{v}_{9}^{\text{RE}}$ -0.00547 -0.00568 -0.00598
$\bm{v}_{10}^{\text{RE}}$ -0.00622 -0.00286 -0.00256
$\bm{v}_{1}^{\text{CAR}}$ -0.000497 0.00202 0.00225
$\bm{v}_{2}^{\text{CAR}}$ 0.000674 -0.00275 -0.00306
$\bm{v}_{3}^{\text{CAR}}$ 0.000991 -0.00424 -0.00449
$\bm{v}_{4}^{\text{CAR}}$ -0.000320 0.00175 0.00145
$\bm{v}_{5}^{\text{CAR}}$ -0.000353 0.00144 0.00160
$\bm{v}_{6}^{\text{CAR}}$ -0.000915 0.00436 0.00415
$\bm{v}_{7}^{\text{CAR}}$ 0.000147 -0.000364 -0.000669
$\bm{v}_{8}^{\text{CAR}}$ 0.0000614 -0.000583 -0.000278
$\bm{v}_{9}^{\text{CAR}}$ -0.000451 0.00174 0.00205
$\bm{v}_{10}^{\text{CAR}}$ 0.000986 -0.004170 -0.004470
$\bm{v}_{1}^{\text{GP}}$ -0.0200 -0.0178 -0.0175
$\bm{v}_{2}^{\text{GP}}$ -0.0140 -0.0117 -0.0114
$\bm{v}_{3}^{\text{GP}}$ 0.00635 0.00377 0.00347
$\bm{v}_{4}^{\text{GP}}$ -0.00901 -0.0150 -0.0148
$\bm{v}_{5}^{\text{GP}}$ -0.0115 -0.00942 -0.00914
$\bm{v}_{6}^{\text{GP}}$ 0.00187 0.00787 0.00818
$\bm{v}_{7}^{\text{GP}}$ -0.00104 0.000352 0.000657
$\bm{v}_{8}^{\text{GP}}$ 0.000827 0.000184 -0.000120
$\bm{v}_{9}^{\text{GP}}$ 0.00366 0.00838 0.00869
$\bm{v}_{10}^{\text{GP}}$ -0.00301 -0.00223 -0.00254
(b) Balance and representativeness: latent covariates.